# A Comprehensive Survey on Data Augmentation

### v3

@@ IV-B5 Sequence Transformation
Unlike Value Perturbation, sequence perturbation is implemented on the X-axis, either on the time domain or the frequency domain. Window Slicing splits the time-series data into pieces. It is also the basis of many Sequence Morphing methods because they apply varied operations on these slices.
@@
Inspired by image cropping, which is widely used for image data augmentation, MCNN [109] proposes window slicing. It extracts slices from the entire sequence as new training samples. Each slice is assigned with the label of the original sequence.
@@
Based on Window Slicing, [110] extends or compresses the generated spans, which is called window wrapping. This is equivalent to speeding up or slowing down a span of the sequence. Vowel stretching [111] is a special kind of window wrapping technique. It only perturbs the sequence by prolonging the vowels. This is meant to be consistent with how children speak. [67] propose to randomly permeate these slices and concatenate them as a long sequence. Similar processing is also called Random Shuffling and is used in [112]. VTLP [113] applies window wrapping in the frequency domain. It adds a small distortion to the central frequency. SpecAugment [114] also proposes to augment the frequency domain. It generates masks of length sampled from a uniform distribution and uses each of them to mask consecutive frequency channels.
@@ IV-C1 Image Value-structure Transformation
Policy-based methods search for the optimized augmentation policy that includes both value-based transformations and structure-based transformations, as demonstrated in Fig. 2f. Prompt-based Image Editing employs a text-to-image model to change the content, background, or style of the image as humans desire. When implementing the augmentation, both coloring and the positions of pixels are changed (Fig. 2g).
@@
AutoAugment [13] represents data augmentation as a decision-making process and uses reinforcement learning to search for the best operations. In AutoAugment, a policy includes a group of basic data augmentation operations with probabilities and magnitudes. AutoAugment processes images under the policy, evaluates these augmented images on downstream networks, and rewards a policy generator for better performance. Fast AutoAugment [115] proposes a more efficient searching algorithm for augmentation policies, which works significantly faster than AutoAugment and thus allows a larger number of sub-policies, which should benefit the generalization performance. RandAugment [116] removes the need for a separate search phase on a proxy task, and makes the training process less complicated and less expensive. Population Based Augmentation [117] learns an augmentation schedule rather than a fixed policy as in AutoAugment to improve efficiency. These methods combine the benefits of both value-based transformations and structure-based transformations and further discover the potential of data augmentation with individual image data.
@@
Visual generative models have been rapidly evolving in recent years. These models are typically conditioned on text embeddings, allowing flexible image editing by modifying textual descriptions of images and prompting the models. DA-Fusion [118] learns fine-grained concepts, replaces the original category labels, and uses the Stable Diffusion model to generate new images. SaSPA [119] uses GPT-4 to generate prompts based on the sample’s meta-class, then inputs the prompts, reference images, and detected edges to the Diffusion model to generate augmented images.
@@ IV-C2 Text Value-structure Manipulation
Hierarchical Data Augmentation explicitly augments both value and structure information of text data. Back-translation Augmentation and Paraphrase Augmentation use neural networks to generate sentences that are semantic-preserving but different from original sentences in both wording and syntax.
@@
Hierarchical Data Augmentation (HDA) [120] augments text data at both word and sentence levels. It evaluates the roles of words and sentences by training a hierarchical attention network to obtain the attention values of the text data at both levels. Attention values help decide the importance of components and augment the data by cropping and concatenating these components.
Back-translation Augmentation
Back-translation Augmentation. Back-translation [121] means translating text data from one language to another, or a sequence of other languages, and finally translating back to the original language. Semantic information is preserved, but the translation process can introduce some variance in syntax and wording. Such transformation can help the model learn more flexible expressions that convey the same meaning.
@@
AugGPT [122], for example, prompts the state-of-the-art ChatGPT model to rewrite sentences in the dataset and preserves dataset coherence after data augmentation. Taboo [123] requires paraphrasing without using some frequently used words, Chain [124] paraphrase one sentence multiple times, and Hint [124] provide some examples, to encourage the LLM response otherwise. DoAug [125] constructs a preference dataset based on the diversity of paraphrased sentences and trains the LLM with the DPO to boost its generation diversity.
@@ V Multi-instance Level Data Augmentation
Multi-instance augmentation combines value and structure information from different samples. Specific operation varies depending on the data modality in which it is applied.
@@ V-A Value-based Mixture
V-A1 Pixel Mixture
Image Mixup computes an arithmetical mixture of the color from different images and Neural Blending resorts to neural networks to enhance such process.
Image Mixup
Mixup [126] mixes two random images on the input layer by calculating their interpolation pixel-wise. Pixels from two original images each contribute a portion to the new image:
$\vspace{-2mm}\mathbf{\tilde{x}}=\lambda\mathbf{x}_{i}+(1-\lambda)\mathbf{x}_{j},\qquad\mathbf{\tilde{y}}=\lambda\mathbf{y}_{i}+(1-\lambda)\mathbf{y}_{j}$ (14)
@@ V-A1 Pixel Mixture
where $(\mathbf{x}_{i},\mathbf{y}_{i})$ and $(\mathbf{x}_{j},\mathbf{y}_{j})$ are two labeled images from the dataset, and $\lambda\in[0,1]$. As a variant, manifold mixup [127] is operated on the hidden layers of the neural network. Most mixup methods combine two different images, while AugMix [128] applies three operations to one image, and combines the three augmented images to generate a new sample. In some approaches, the mixing operation varies in strength across regions. For example, SmoothMix [129] applies a soft-edged mask when mixing. Most mixup methods create new samples with soft labels, as in Eq.14, while there are some exceptions. Sample Paring [130], for example, uses the label of one of the original images, and that of the other image is left unused. ReMix [131] addresses the class imbalance problem by assigning the minority class label to augmented data.
Neural Blending
AutoMix [132] integrates a Mix Block in the training process. It builds a bridge between the selection of mixup policy and the optimization of the model. In AutoMix, the augmentation module is optimized along with the classification feature map, and the mixup functions and mixing ratio will updated according to the classification loss. Another image blending technique is proposed as Smart Augmentation [133]. At the input end of Smart Augmentation, two random samples are fed into the network through two channels, and a CNN sub-network yields a one-channel output with the same shape as the input, which means one new image is generated based on the two input images. The augmented image is then used for downstream tasks, and the augmentation sub-network adjusts to the task performance accordingly.
V-A2 Text Representation Mixture
@@
Unlike image data, whose coloring information is expressed by numerical values and can be interpolated at the input layer, words are difficult to interpolate directly, so text interpolation is usually done at the embedding layer.
Text Mixup
The adaptation of Mixup in sentence classification is studied in [134], which proposes wordMixup and senMixup. WordMixup interpolates a pair of word embeddings and then feeds the interpolation to the sentence encoder. SenMixup, on the other hand, interpolates a pair of sentence embeddings. The embedding vector produced by WordMixup or SenMixup is then passed to the softmax layer to produce the target distribution of the classification task. The use of mixup in text data augmentation has also been extensively studied in many research, such as nonlinear Mixup [135] and DoubleMix [136]. GPT3Mix [137] makes use of the latest large language model to mix sentences. By inserting two source sentences in a prompt template that describes the instruction to the model, we can get a mixed sentence of high quality with a soft label.
V-A3 Node Mixture
Feature or label information from different nodes can be mixed by Graph Propagation, which is often used for tasks such as node classification.
Graph Propagation
@@ V-A3 Node Mixture
Propagation refers to spreading information from one node to other nodes along the graph. These methods include feature propagation and label propagation. Feature propagation shares node features with other nodes. After feature vectors of some nodes are masked out, GRAND [50] further augments the node features by randomly propagating them through edges. MV-GCN [138] first creates three different views from the original graph, propagates the features in each, and then combines the three representations. Label propagation algorithm (LPA) [139] propagates labels from labeled nodes to all unlabeled nodes according to their proximity through areas with dense unlabeled nodes. AutoGRL [140] applies LPA in their semi-supervised framework to label more nodes utilizing a few ground truth nodes. [141] performs two types of label propagation, correcting the base predictions and smoothing the final prediction, respectively. Though not in a typical graph-based scenario, [142] constructs a nearest neighbor graph on the classifier’s manifolds. The affinity matrix, analogous to the adjacency matrix in this case, is constructed based on the $k$ nearest neighbors of each node’s embeddings, and this matrix decides how labels propagate.
V-A4 Cell Value Mixture
Due to its well-aligned structure, tabular data is very suitable for interpolation.
Tabular Mixup
@@ V-A4 Cell Value Mixture
SMOTE [37] is among the first works to study data interpolation, even before the proposal of Mixup [126]. It addresses the problem of class imbalance by oversampling from the minority class. The samples are selected within the $k$ nearest neighbors in the feature space and new synthetic data are produced by a random interpolation operation. Since tabular features are often nominal, two variants, namely SMOTE-NC and SMOTE-N, are also proposed along with the initial version of SMOTE [37]. As argued in [143], SMOTE is subject to unrealistic samples when interpolating in a sparse feature space. To solve this problem, TAEI [143] first encodes the sparse data into a dense latent space, interpolates the new data, and then decodes them to obtain the original data plus the augmented data. Contrastive Mixup [144] applies mixup for tabular data in a self- and semi-supervised learning framework. SAINT [145] combines CutMix [146] with contrastive learning to learn from tabular data. ExcelFormer [147] proposes two schemes, namely Feat-Mix and Hidden-Mix. They interpolate data in the input layer and the hidden layer, respectively.
V-A5 Sequence Value Mixing
@@
Methods that combine information from different sequences are referred to as sequence averaging or sequence mixing. DTW Barycenter Averaging (DBA) [148] revolves around the averaging problem for Dynamic Time Warping. It develops a global technique for computing the average of a set of sequences and avoids using iterative pairwise averaging. In Weighted-DBA [149], instead of each time series contributing equally to the final average, some can contribute more than others, so that an infinite number of new examples from any set of given time series can be generated. [150] applies both interpolation and extrapolation in the feature space to augment time-series data. It first uses a sequence autoencoder to learn a feature space from unlabeled data. The data is then encoded and augmented with additive noises, interpolation, and extrapolation. Finally, the data in feature spaces is decoded and classified. Extending sequence mixing to the frequency domain, [151] applies the Equalized Mixture Data Augmentation (EMDA) to create new data. It computes the weighted average of two randomly chosen spectrograms with the same label.
@@ V-B Structure-based Combination
V-B1 Image Structural Combination
Image Patching puts together small fragments of many images or puts a part of an image over a second image.
Image Patching
RICAP [152] randomly crops four images, and patches them to construct a new image. CutMix [146] makes use of two images. It is developed based on Cutout [9], but the removed region is replaced by a patch from another image. Attentive CutMix [153] further enhances the CutMix method by adding a feature extraction module that selects an important or representative region, and pastes this attentive patch onto the other image. CowMask [154] introduces patching with irregular shape rather than rectangular.
V-B2 Text Structural Combination
Text Fragment Merging combines text data by merging text fragments from one sentence into another.
Text Fragment Merging
GECA [155] aims to remove a fragment from a sentence and populate it with another fragment that occurs in some common environment. Subtree Swapping [156], Substructure Substitution [157], and TreeMix [158] use tree structures to analyze the part of speech or role in the sentences and thus make reasonable swaps. SSMix [159] replaces a low saliency segment in a sentence with a segment with high saliency from another sentence and mixes sentence labels according to the ratio of their span length in the combined sentence.
@@ V-B3 Table Concatenation
Contrary to single-instance level tabular augmentation, which lacks structural information to use, multi-instance level tabular augmentation suffers from strict and fixed table structures. Columns differ in their nomenclatures and representations, so it is hard to combine and learn from multiple distinct tables. If these restrictions are relaxed, models can learn across different tables, be more transferable, and better predict unseen tables.
Structure Relaxation
TransTab [160] suggests that multiple tables could share partially overlapped columns in the real world, but traditional methods often fail to fully utilize the data due to the removal of non-overlapping columns and mismatched samples when cleaning the data. TransTab first divides the table into three parts, each containing categorical, binary, and numerical data. These three types of data undergo different operations, are all converted to embeddings, and are again integrated into a new and encoded table. The processed tables can be joined in many ways, depending on the application scenarios. The augmented table contains information from many original tables and helps the model learn across them.
@@ V-C1 Image Value-structure Mixture
Diffusion Models take textual prompts as inputs and can perform a flexible and smart mixture of existing images.
@@
V-C2 Graph Value-structure Mixutre
Graph Mixup refers to mixing graph nodes or graph instances. This often involves both node attribute values and the topology structures.
Graph Mixup
Due to the irregular connectivity and various topological structures of graph data, it is hard to directly combine nodes or graphs. For node-level mixup, one solution to this problem is temporarily ignoring the edges and mixing node features in latent space. GraphMix [163] interpolates the hidden states and labels of input nodes. A topology-involved node mixup method is proposed in [164]. It randomly pairs nodes and includes the nodes’ local topology representations when interpolating them. S-Mixup (Structural Mixup) [165] considers neighboring edges when pairing nodes but works in the input space and discusses how to connect the node with the original graph with edges. It matches nodes based on GNN classification confidence and connects with high-gradient edges to the mixed node. For graph-level mixup, ifMixup [166] adds dummy nodes so that two graphs have the same number of nodes. These dummy nodes have all-zero feature vectors and are not connected to any other nodes, or in other words, the weights of edges on dummy nodes are zeros. S-Mixup (Graph Mixup with Soft Alignments) [167] also aligns graphs first before mixing them. It obtains an assignment matrix by a soft assignment and uses the assignment matrix to transform a graph so that every node corresponds with a node in the other graph. Compared with the above approaches, G-Mixup [168] proposes a more sophisticated way to interpolate graph data. It first uses graphs within the same class to estimate a graphon, i.e. graph generator:
$\vspace{-2mm}\mathbf{\mathcal{G}}\rightarrow W_{\mathbf{\mathcal{G}}},\mathbf{\mathcal{H}}\rightarrow W_{\mathbf{\mathcal{H}}}$ (15)
The graphon $W:[0,1]^{2}\rightarrow[0,1]$ is a matrix, where $W(i,j)$ describes the probability of an edge between nodes $i$ and $j$, and graphs of the same class are considered to be generated from the same graphon. Graphons of two classes of graphs are then interpolated to produce a mixed graphon:
$\vspace{-2mm}W_{\mathbf{\mathcal{I}}}=\lambda W_{\mathbf{\mathcal{G}}}+(1-\lambda)W_{\mathbf{\mathcal{H}}}$ (16)
By sampling from the graphon mixup $W_{\mathbf{\mathcal{I}}}$, we acquire a series of graphs with the same soft labels:
$\vspace{-2mm}\mathbf{y}_{\mathbf{\mathcal{I}}}=\lambda\mathbf{y}_{\mathbf{\mathcal{G}}}+(1-\lambda)\mathbf{y}_{\mathbf{\mathcal{H}}}$ (17)
During graphon estimation, G-Mixup aligns the original node features and obtains the graphon node features average pooling, which are used as node features of the generated graphs. DAGAD [169] mixes graph representations by performing random permutations, demonstrating that data augmentation is helpful in scenarios such as graph anomaly detection.

### v4

@@ IV-B5 Sequence Transformation
Unlike value perturbation, sequence perturbation is implemented on the X-axis. Window slicing splits the time-series data into pieces. It is also the basis of many sequence morphing methods because they apply varied operations on these slices.
@@
Inspired by image cropping, MCNN [109] proposes window slicing. It extracts slices from the entire sequence as new training samples. Each slice is assigned the label of the original sequence.
@@
Based on window slicing, [110] extends or compresses the generated spans, which is called window wrapping. This is equivalent to speeding up or slowing down a span of the sequence. Vowel stretching [111] is a special kind of window wrapping technique. It only perturbs the sequence by prolonging the vowels. This is meant to be consistent with how children speak. Random shuffling [67, 112], randomly permutates these slices and concatenates them as a long sequence. VTLP [113] applies window wrapping in the frequency domain. It adds a small distortion to the central frequency. SpecAugment [114] also augments the frequency domain. It masks consecutive frequency channels of length sampled from a uniform distribution.
@@ IV-C1 Image Value-structure Transformation
Policy-based methods search for the optimized augmentation policy that includes both value-based and structure-based transformations, as demonstrated in Fig. 2f. Prompt-based image editing employs a text-to-image model to change the content, background, or style of the image as humans desire, changing both the coloring and the position of pixels (Fig. 2g).
@@
A policy includes a sequential group of basic augmentation operations with probabilities and magnitudes. AutoAugment [14] represents data augmentation as a decision-making process and uses reinforcement learning to search for the best policy. It processes images under the policy, evaluates these augmented images on downstream networks, and rewards a policy generator for better performance. Fast AutoAugment [115] proposes a more efficient searching algorithm, which is faster and thus allows a larger number of sub-policies, benefiting the generalization performance. These methods rely on independent reinforcement learning processes for each task, and the searched policies do not guarantee optimality when transferred to other tasks, making the methods cost-inefficient. RandAugment [116] removes the need for a separate search phase on a proxy task and makes the training process less complicated and less expensive. Population Based Augmentation [117] learns an augmentation schedule rather than a fixed policy as in AutoAugment to improve efficiency. These methods combine the benefits of both value-based transformations and structure-based transformations and further discover the potential of data augmentation with individual image data.
@@
Visual generative models are typically conditioned on text embeddings, allowing flexible image editing by modifying textual descriptions of images and prompting the models. DA-Fusion [118] learns fine-grained concepts, replaces the original category labels, and uses the Stable Diffusion model to generate new images. SaSPA [119] uses GPT-4 to generate prompts based on the sample’s meta-class, then inputs the prompts, reference images, and detected edges to the diffusion model to generate augmented images. Generating the final augmentation set with diffusion models can be labor-intensive because models are very sensitive to the prompts, so users must prepare detailed image descriptions carefully and might have to examine the prompts repeatedly.
@@ IV-C2 Text Value-structure Manipulation
Hierarchical data augmentation explicitly augments both value and structure information of text data. Back-translation and paraphrase augmentation use neural networks to generate sentences that are semantic-preserving but different from original sentences in both wording and syntax.
@@
Hierarchical data augmentation [120] augments text data at both word and sentence levels. It evaluates the roles of words and sentences by training a hierarchical attention network to obtain the attention values of the text data at both levels. Attention values help decide the importance of components and augment the data by cropping and concatenating these components.
Back-translation
[121] proposes to translate text data from one language to another, or a sequence of other languages, and finally translates back to the original language. Semantic information is preserved, but the translation process can introduce some variance in syntax and wording. Such transformations can help the model learn more flexible expressions that convey the same meaning.
@@
AugGPT [122], for example, prompts the state-of-the-art large language model (LLM) to rewrite sentences in the dataset and preserves dataset coherence after data augmentation. Taboo [123] requires paraphrasing without using some frequently used words, Chain [123] paraphrases one sentence multiple times, and Hint [123] provides some examples to encourage the LLM response otherwise. DoAug [124] constructs a preference dataset based on the diversity of paraphrased sentences and trains the LLM with the DPO to boost its generation diversity. Although such paraphrase methodology is capable of preserving sentence meanings, it is still possible that the paraphraser misunderstands the nuanced semantics and causes semantic drift that invalidates the sample’s label.
@@ V Multi-instance Level Data Augmentation
Multi-instance augmentation combines value and structure information from different samples. The specific operation varies depending on the data modality in which it is applied.
@@ V-A Value-based Mixture
V-A1 Image Mixup
@@
Mixup [125] mixes two random images on the input layer by calculating their interpolation pixel-wise. Pixels from two original images each contribute a portion to the new image: $\mathbf{\tilde{x}}=\lambda\mathbf{x}_{i}+(1-\lambda)\mathbf{x}_{j},\mathbf{\tilde{y}}=\lambda\mathbf{y}_{i}+(1-\lambda)\mathbf{y}_{j}$, where $(\mathbf{x}_{i},\mathbf{y}_{i})$ and $(\mathbf{x}_{j},\mathbf{y}_{j})$ are two labeled images from the dataset, and $\lambda\in[0,1]$. Manifold mixup [126] is operated on the hidden layers of the neural network. AugMix [127] applies three operations to one image and combines the three augmented images to generate a new sample. SmoothMix [128] applies a soft-edged mask to mix different regions in varied strength. Most mixup methods create new samples with soft labels, while sample pairing [129] uses the label of one of the original images, and that of the other image is left unused. Similarly, ReMix [130] addresses the class imbalance problem by assigning the minority class label to augmented data.
V-A2 Text Mixup
@@
Unlike image data, whose coloring information is expressed by numerical values and can be interpolated at the input layer, words are difficult to interpolate directly, so text interpolation is usually done at the embedding layer. The adaptation of Mixup in sentence classification is studied in [131], which proposes wordMixup and senMixup. Its wordMixup interpolates a pair of word embeddings and then feeds the interpolation to the sentence encoder. On the other hand, senMixup interpolates a pair of sentence embeddings. The use of mixup in text data augmentation has also been studied in other research, such as nonlinear Mixup [132] and DoubleMix [133].
V-A3 Graph Propagation
@@
Propagation refers to spreading information from one node to other nodes along the graph. These methods include feature propagation and label propagation. Feature propagation shares node features with other nodes. After feature vectors of some nodes are masked out, GRAND [50] further augments the node features by randomly propagating them through edges. MV-GCN [134] first creates three different views from the original graph, propagates the features in each, and then combines the three representations. The label propagation algorithm (LPA) [135] propagates labels from labeled nodes to all unlabeled nodes according to their proximity through areas with dense unlabeled nodes. AutoGRL [136] applies LPA in its framework to label more nodes utilizing a few ground truth nodes. [137] performs two types of label propagation, correcting the base predictions and smoothing the final prediction, respectively.
V-A4 Tabular Mixup
@@
Due to its well-aligned structure, tabular data is very suitable for interpolation. SMOTE [37] is among the first works to study data interpolation. It addresses the problem of class imbalance by oversampling from the minority class. The samples are selected from the $k$ nearest neighbors in the feature space, and new synthetic data are produced by a random interpolation operation. As argued in [138], SMOTE is subject to unrealistic samples when interpolating in a sparse feature space. To solve this problem, TAEI [138] first encodes the sparse data into a dense latent space, interpolates the new data, and then decodes them to obtain the original data plus the augmented data. Contrastive Mixup [139] applies mixup for tabular data in a self- and semi-supervised learning framework. SAINT [140] combines CutMix [141] with contrastive learning to learn from tabular data. ExcelFormer [142] proposes two schemes, namely Feat-Mix and Hidden-Mix. They interpolate data in the input layer and the hidden layer, respectively.
V-A5 Sequence Mixing
@@
DTW Barycenter Averaging (DBA) [143] revolves around the averaging problem for Dynamic Time Warping. It develops a global technique for computing the average of a set of sequences and avoids using iterative pairwise averaging. In Weighted-DBA [144], instead of each time series contributing equally to the final average, some can contribute more than others, so that an infinite number of new examples from any set of given time series can be generated. [145] applies both interpolation and extrapolation in the feature space to augment time-series data. It first uses a sequence autoencoder to learn a feature space from unlabeled data. The data is then encoded and augmented with additive noise, interpolation, and extrapolation. Finally, the data in feature spaces is decoded and classified. Extending sequence mixing to the frequency domain, [146] applies the Equalized Mixture Data Augmentation (EMDA) to create new data. It computes the weighted average of two randomly chosen spectrograms with the same label.
@@ V-B Structure-based Combination
V-B1 Image patching
Image patching puts together small fragments of many images or puts a part of an image over a second image. RICAP [147] randomly crops four images and patches them to construct a new image. CutMix [141] is developed based on Cutout [10], but the removed region is replaced by a patch from another image. Attentive CutMix [148] further enhances the CutMix method by adding a feature extraction module that selects an important or representative region and pastes this attentive patch onto the other image. CowMask [149] introduces patching with an irregular shape rather than a rectangular one.
V-B2 Text Fragment Merging
Text fragment merging combines text data by merging text fragments from one sentence into another. GECA [150] removes a fragment from a sentence and populates it with another fragment that occurs in some common environment. Subtree Swapping [151], Substructure Substitution [152], and TreeMix [153] use tree structures to analyze the part of speech or role in the sentences and thus make reasonable swaps. SSMix [154] replaces a low-saliency segment in a sentence with a segment with high saliency from another sentence and mixes sentence labels according to the ratio of their span length in the combined sentence.
@@ V-B3 Table Concatenation
Contrary to single-instance level tabular augmentation, which lacks structural information to use, multi-instance level tabular augmentation suffers from strict and fixed table structures. Columns differ in their nomenclatures and representations, so it is hard to combine and learn from multiple distinct tables. If these restrictions are relaxed, models can learn across different tables, be more transferable, and better predict unseen tables. TransTab [155] suggests that multiple tables could share partially overlapped columns in the real world, but traditional methods often fail to fully utilize the data due to the removal of non-overlapping columns and mismatched samples when cleaning the data. TransTab first divides the table into three parts, each containing categorical, binary, and numerical data. These three types of data undergo different operations, are all converted to embeddings, and are again integrated into a new and encoded table. The processed tables can be joined in many ways, depending on the application scenarios. The augmented table contains information from many original tables and helps the model learn across them.
@@ V-C1 Image Value-structure Mixture
CNN models merge images by fusing pixels through channels and across the region within the reception fields. Diffusion models take textual prompts as inputs and can perform a flexible and smart mixture of existing images.
Neural Blending
AutoMix [156] integrates a mix block in the training process. It builds a bridge between the selection of the mixup policy and the optimization of the model. Another image blending technique is proposed as Smart Augmentation [157]. At its input end, two random samples are fed into the network through two channels, and a CNN sub-network yields a one-channel output with the same shape as the input, which means one new image is generated based on the two input images. The augmented image is then used for downstream tasks, and the augmentation sub-network adjusts to the task performance accordingly.
@@
V-C2 Text Value-structure Mixture by Generative Mixing
Generative mixing leverages the text understanding and generation capabilities of LLMs to combine information from multiple sentences. GPT3Mix [160] makes use of the latest large language model to mix sentences. By inserting two source sentences in a prompt template that describes the instruction to the model, we can get a mixed sentence of high quality with a soft label. PromptMix [161] also prompts the LLM to combine information from two sentences and then uses another LLM to relabel the generated augmentation sentence. Combining information from two text pieces is a subtle problem, and the specific behaviors of the LLMs highly depend on the prompts, which include the instruction, task description, and few-shot examples. So, the users have to try hard to design an ideal prompt so that the model works well.
V-C3 Graph Value-structure Mixture by Graph Mixup
Graph mixup refers to mixing graph nodes or graph instances. This often involves both node attribute values and the topology structures. For node-level mixup, a topology-involved method [162], randomly pairs nodes and includes the nodes’ local topology representations when interpolating them. Structural Mixup [163] considers neighboring edges when pairing nodes and discusses how to connect the node with the original graph with edges. For graph-level mixup, ifMixup [164] adds dummy nodes so that two graphs have the same number of nodes. Graph Mixup with Soft Alignments [165] obtains an assignment matrix by a soft assignment and uses that matrix to transform a graph so that every node corresponds with a node in the other graph. G-Mixup [166] first uses graphs within the same class to estimate a graphon, i.e., graph generator: $\mathbf{\mathcal{G}}\rightarrow W_{\mathbf{\mathcal{G}}},\mathbf{\mathcal{H}}\rightarrow W_{\mathbf{\mathcal{H}}}$. The graphon describes the probability of an edge between nodes $i$ and $j$. Graphons of two classes of graphs are then interpolated to produce a mixed graphon: $W_{\mathbf{\mathcal{I}}}=\lambda W_{\mathbf{\mathcal{G}}}+(1-\lambda)W_{\mathbf{\mathcal{H}}}$. By sampling from the graphon mixup $W_{\mathbf{\mathcal{I}}}$, we acquire graphs with the same soft labels: $\mathbf{y}_{\mathbf{\mathcal{I}}}=\lambda\mathbf{y}_{\mathbf{\mathcal{G}}}+(1-\lambda)\mathbf{y}_{\mathbf{\mathcal{H}}}$. DAGAD [167] mixes graph representations by performing random permutations, demonstrating that data augmentation is helpful in scenarios such as graph anomaly detection.